OpenStudy (anonymous):

Maximum and minimum values of the Functions f(x)=(x^2-4)^2

OpenStudy (zbrandz23):

Cute pic

OpenStudy (anonymous):

minimum when x=2 or -2 f(2)=(2^2-4)^2=(4-4)^2=0 f(-2)=((-2)^2-4)^2=0 max + and - infinite